OpenStudy (anonymous):

Determine whether y varies directly with x. If so, find the constant of variation k. 2y = –7x – 23

OpenStudy (anonymous):

y=-7/2x-23/2 k=-7/2

OpenStudy (anonymous):

so y does not varie directly with x?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

Ok